Te value, v, of a shopkeeper's inventory software program is given by v = -1/2t + 3, where t is the number of years the shopkeeper bought the program. What is the program worth 6 years after it is purchased? Plot the point with the given t-value, 6.
Answer by user_dude2008(1862)

v = -1/2t + 3


v = -1/2(6) + 3


v = -6/2 + 3

v = -3 + 3
v=0

Answer: value is $0